OpenStudy (anonymous):

Find each percent of change. Round to the nearest whole percent if necessary. State whether the percent of change is an increase or decrease. original: $20.95 new: $15.10 a. 28% increase b. 28% decrease c. 39% increase d. 39% decrease

OpenStudy (anonymous):

@sky577 @grabill

OpenStudy (anonymous):

a becuase like the other prob you divivde 20.95 by 15.10 and get 1.39 the multiply it by 20.95 and get 29.1 so its a decrease to the new price

OpenStudy (bookworm14):

I got B , I did 20.95 * 39% = 8.1 , then I did 20.95*28% = 5.86. And when you subtract you get 20.95-15.1 = 5.85 , therefore it was A decrease in price of 28%- hope this helps - Bree :)